OpenStudy (anonymous):

A student is trying to solve the system of two equations given below: Equation P: y + z = 6 Equation Q: 3y + 4z = 1 Which of the following is a possible step used in eliminating the y-term?

OpenStudy (anonymous):

@ LegendarySadist

OpenStudy (anonymous):

a. (y + z = 6) ⋅ 4 b. (3y + 4z = 1) ⋅ 4 c. (y + z = 6) ⋅ −3 d. (3y + 4z = 1) ⋅ 3

OpenStudy (anonymous):

please help will give medal

OpenStudy (jameshorton):

you would multiply the first equation by a -3 this will give you : -3y - 3z = -18 3y + 4z = 1 --------------- 0 + z = -17 z = -17 as you can see, when added, the y terms cancel out...they are eliminated answer is C

OpenStudy (anonymous):

i don't understand how you did that?

OpenStudy (jameshorton):

easy

OpenStudy (anonymous):

what do you mean easy?

OpenStudy (jameshorton):

for the first equation multiple A -3

OpenStudy (jameshorton):

what is a -3

OpenStudy (anonymous):

1?

OpenStudy (jameshorton):

do you know what the elimination method is, have you ever done elimination previously?

OpenStudy (anonymous):

no

OpenStudy (jameshorton):

watch this video and tell me what your learned

OpenStudy (jameshorton):

https://www.youtube.com/watch?v=tphSgpilNv8

OpenStudy (jameshorton):

when your done tell me and im ask you a question

OpenStudy (jameshorton):

Well if we want to eliminate the y-term, we need to either make the y term above it equal to it and then subtract (substitution method) or make the number above it equal to it's negative and then add (addition method).

OpenStudy (anonymous):

So if there is an equation like x+y=-3 and x-y=1 you add through columns to find the answer

OpenStudy (jameshorton):

Let's say we have the equations 2x+y=4 4x+3y=7 If we want to get the y alone, we would mutiply the entire top by 2 and then subtract. 2(2x+y=4)→4x+2y=8 With this we would subtract the bottom equation by the top. So 4x+2y=8 −(4x+3y=7) to get −y=1 In this example I was eliminating the x variable with the substitution method. Understand how the process works?

OpenStudy (anonymous):

yes thank you so much!

OpenStudy (jameshorton):

welcme i have to go so bye

OpenStudy (anonymous):

ok bye